# Measurement-induced phase transition in periodically driven free-fermionic systems

### v2

@@ I Introduction
Another class of entanglement phase transitions arises from the continuous monitoring of quantum systems. The stochastic Schrödinger equation typically governs these open quantum dynamics, and averaging over different realizations of the states leads to the well-known Lindblad master equation dynamics Paule (2018). Since entanglement entropy is a nonlinear function of the state, calculating it for each trajectory separately and then averaging reveals a phase transition. In contrast, any linear function of the state is featureless and does not exhibit such transitions. Several studies have explored these entanglement phase transitions in both interacting and non-interacting systems, considering clean as well as disordered systems with long-range and short-range hopping, using both Lindblad jump operators and quantum state diffusion approaches Alberton et al. (2021); Szyniszewski et al. (2023); Cao et al. (2019); Minato et al. (2022a); Müller et al. (2022); Yamamoto and Hamazaki (2023); Fuji and Ashida (2020); Ladewig et al. (2022). Continuously monitored open systems display competition between unitary evolution, which drives entanglement growth, and measurements, which tend to localize the wavefunction by collapsing it into eigenstates of the measurement operators. As a result, these systems undergo a measurement-induced phase transition as a function of measurement strength, which also has been reported recently in experiments Koh et al. (2023a). It has been shown that, in both clean and disordered free-fermionic systems, there is an entanglement phase transition from a critical to an area law phase as the strength of continuous measurements is tuned. This transition was predicted to fall into the Berezinskii-Kosterlitz-Thouless universality class Cao et al. (2019); Szyniszewski et al. (2023); Alberton et al. (2021), a result that was also supported by a theoretical study in which the dynamics of a continuously monitored free fermionic system were described by an effective non-Hermitian sine-Gordon model Buchhold et al. (2021). However, it is important to point out that the stability of the critical phase in a one-dimensional free-fermionic system under random projective measurements has been questioned recently, and it has been argued in the thermodynamic limit, for any finite measurement strength, only the area-law phase will survive Poboiko et al. (2023); Starchl et al. (2024a); Poboiko et al. (2025). Also in the case of Dirac fermions, the same conclusion is claimed recentlyMüller et al. (2025) after refining their previous resultsBuchhold et al. (2021). There are recent theoretical studies of a monitored chain of a system of size $L$ of Majorana fermions, which shows ($\ln L)^{2}$ behavior in the critical phase Fava et al. (2023), in contracts for noninteracting systems where the $U(1)$ symmetry is preserved, reported to show area-law phase for all measurement strength in presence of complex random hoppingFava et al. (2024). On the other hand, in the case of interacting systems, a transition from a volume law to an area law has been observed Fuji and Ashida (2020); Yamamoto and Hamazaki (2023); Xing et al. (2024). Furthermore, the role of measurements in localizing quantum states in nonunitary quantum circuits has also attracted significant interest Zabalo et al. (2022); Sierant and Turkeshi (2022); Oshima and Fuji (2023); Weinstein et al. (2023); Côté and Kourtis (2022); Block et al. (2022); Sang and Hsieh (2021); Lunt et al. (2021); Li et al. (2021); Fan et al. (2021); Zabalo et al. (2020); Turkeshi et al. (2020); Gullans and Huse (2020a, b); Bao et al. (2020); Choi et al. (2020); Tang and Zhu (2020); Jian et al. (2020). Some of these studies have also been extended in dimension $D>1$  Chahine and Buchhold (2024); Jin and Martin (2024); Poboiko et al. (2024) and a semi-classical limit of such transition has also been introduced recently Ruidas and Banerjee (2024). In contrast to disorder-induced localization observed in the Anderson-localized or MBL phase, the measurement-induced localization is governed by a completely different mechanism and driven by quantum measurements and measurement strength.
@@
In parallel, in the last few decades, periodically driven (Floquet) quantum systems have emerged as a powerful framework for studying non-equilibrium phases of matter Banerjee and Sengupta (2024); Sarkar et al. (2021, 2022); Nag et al. (2014); Agarwala et al. (2016); Luitz et al. (2018); Chakrabarty and Datta (2024a). Floquet systems are known to exhibit exotic phases, such as discrete-time crystals Khemani et al. (2016); Else et al. (2016); Zhang et al. (2017) and Floquet topological insulators Oka and Aoki (2009); Lindner et al. (2011); Kitagawa et al. (2011); Kundu et al. (2014); Mukherjee et al. (2016). The fate of many-body localization (MBL) in the presence of a periodic drive has also been studied Lazarides et al. (2015); Ponte et al. (2015). While it has been demonstrated that Floquet systems can stabilize or even induce new phases, it remains an open question how continuous monitoring affects these driven systems and whether the introduction of a periodic drive can significantly alter the dynamics of a continuously monitored system, either enhancing or suppressing the effects of measurements. The effect of time-dependent Hamiltonian has been studied within the Lindblad master equation formalism Scopa et al. (2019); Ikeda et al. (2021), but the effect of periodic drive on a continuously monitored open quantum system, particularly in terms of trajectory-averaged entanglement entropy, remains relatively unexploredLi et al. (2019). In our study, we investigate such cases, and our main finding, based on our accessible system size numerics, is that periodically driven non-interacting systems tend to exhibit a signature of undergoing a BKT transition for reasonably high driving periods. In general, for a given drive protocol, the low-frequency regime favors the critical phase more, implying that if the BKT transition persists, then the critical measurement strength (transition point) increases with the time period. Moreover, the symmetric drive of the hopping amplitude about a zero mean is special; such a system always favors the area-law phase in the thermodynamic limit, independent of the frequency regime, in contrast to what has been observed in the case of the disorder-induced scenario Sarkar et al. (2021). We summarize our setup and main results based on the finite-sized numerical data in the form of a schematic in Fig. 1, which includes real data. We plot the maximum of the local effective central charge, which is a measure of the maximum growth rate of the steady state entanglement with sub-system size (defined later in the manuscript in Eq. (19)) with time-period $T$ for different measurement strength $\gamma$. Note that there is a growing belief based on recent analytical studies that in the no-drive scenario, any infinitesimal measurement strength is enough to stabilize the area law entanglement phase in the thermodynamic limit, but clear numerical evidence is still absent, specifically for the measurement protocols used in this manuscript. Moreover, to the best of our knowledge, no analytical prediction is available for driven systems till now. This work aims to do a detailed comparative numerical analysis between the driven system with a reasonably low-frequency regime (in the main text) and the no-drive/high-frequency regime (in the Appendix D) results for a finite-sized system.
@@ IV.1 High-frequency regime: Analytical prediction
Note that this analysis has been questioned recently with the growing belief of the absence of measurement-induced phase transition in these models, and the analysis predicts a critical to area law phase transition in the no-drive scenario. However, since this study still captures the essential features of the measurement-induced transition and can be extended to the driven system—at least in the extreme high-frequency limit—it motivates us to carry out this analysis to develop some intuition about the driven case.
@@
It is straightforward to check that for a fixed non-zero $\gamma$, in the limit $\epsilon\to 0$, this correlation tends to zero, and the effective central charge $c(\gamma)$ (note that the leading order term in the entanglement entropy $S=\frac{c(\gamma)}{3}\ln l$) will also approach zero as $\sqrt{\frac{\epsilon}{\gamma}}$. This implies that in the case of $\epsilon=0$, for any finite $\gamma$, one would expect to observe the area-law phase. In a nutshell, we make four main predictions based on this RG analysis about the high-frequency limit drive: 1) the entanglement dynamics can be mimicked by the dynamics of no-drive continuously monitored free-fermionic chain with the hopping amplitude $J_{eff}=\epsilon/2$, 2) If one starts in critical-phase in the no-drive regime, i.e. $\epsilon=2J$, with decreasing $\epsilon$, one can be able to probe critical to area-law phase transition, 3) $\epsilon=0$ limit will not display entanglement transition, it will always show the area-law entanglement for any non-zero finite measurement strength, 4) the transition point $\gamma_{c}$ is expected to decay to zero linearly with decreasing $\epsilon$ in the high-frequency region.
In this context, we would like to mention that very recently, a modified version of the RG analysis on sine-Gordon type models has been proposed after refining their previous analysis, in Ref. Müller et al. (2025), which shows the critical to area law transition point shifts towards zero in the case of a non-interacting monitored system in the absence of drive. However, the symmetries and measurement protocols of these continuum models are not exactly identical to our lattice setup. In monitored systems, energy conservation is broken, and measurements can populate fermions across the entire band. Hence, there is no guarantee that any such models (that include Hamiltonian Eq. (9) too) can describe the dynamics of monitored fermions on a tight-binding lattice. Also, the prediction of Ref. Müller et al. (2025) of a logarithmic to area law phase transition in the interacting case, contradicts the other recent works Poboiko et al. (2025); Xing et al. (2024); Guo et al. (2024), where they observed a volume-law to area-law phase transition with increasing measurement rate. Overall, it implies that it is a topic of debate what is the appropriate effective continuum model and the corresponding correct RG equations that can describe the high-frequency limit of a driven continuously monitored system, but in our study, it helps us to build an intuition which we are going to test in the subsequent section with the finite-sized numerics.
@@ IV.2 Numerical results
As a beginning, we study the time evolution of the entanglement entropy, i.e., averaged over many different trajectories, starting from the initial zero entanglement product state for different values of $\epsilon$ and measurement strengths $\gamma,$ for $T=5$. Figure. 6 shows the dynamics of $\overline{S(nT)}$ for $\epsilon=3J/4$. The upper panel shows that for $\gamma=0.05$, the saturation value of the entanglement entropy increases with system size; on the other hand, the lower panel shows that for $\gamma=0.2$, the steady state value of the average entanglement entropy hardly changes with system size. This is a signature that, with increasing measurement strength, an entanglement transition may occur, where for small values of $\gamma$, the steady-state entanglement increases with $L$, and for large $\gamma$, it does not scale with $L$. The inset of the upper panel shows the entanglement dynamics in the high-frequency regime $T=0.5$ for $\gamma=0.05$ and $L=250$ (red-dotted line). As predicted in the previous section, we find that, indeed, for $T=0.5$, the entanglement dynamics can be mimicked by the continuously monitored no-drive scenario in a tight-binding lattice with an effective hopping strength $J_{eff}=\epsilon/2$, which is denoted by points in the inset figure. Given, the results for $T<1$ are almost indistinguishable from the no-drive results, at least for our accessible system size, hence, we consider $T=5>1$ for most of our numerics. The no-drive results are separately presented in the Appendix. D.
@@
where, $\tilde{l}=\frac{L}{\pi}\sin(\pi l/L)$ is the cord length and $l$ is the size of the subsystem varies from $1$ to $L/2$. For logarithmic growth of the entanglement entropy, this measure should be constant; for volume-law scaling, it should scale as $c^{eff}_{l}\sim l$, while area-law behavior leads to $c^{eff}_{l}=0$. To calculate $c^{eff}_{l}$ numerically, we choose $N_{l}$ values of $l_{i}$ with $l_{1}=1$ and $l_{N_{l}}=L/2$ such that the successive ratio of $\frac{\tilde{l^{\prime}}}{\tilde{l}}$ is approximately constant. In Ref. Starchl et al. (2024b), it was shown that the critical phase is bounded from below by $l\sim\gamma^{-1}$ and from above by $l\sim\gamma^{-2}$. If the effective central charge increases with system size and attains a maximum, then it slowly starts decreasing, which indicates a crossover to area law. In Fig. 10 we present the effective central charge $c^{eff}_{l}$ data for the driving period $T=5$. We see a very clear evidence of bending of $c^{eff}_{l}$ after a critical $l^{*}_{c}$ for $\gamma>\gamma_{c}=0.05\pm 0.014$ (black circles represent the maximum). We fit those $l^{*}_{c}$ and obtain the scaling $l^{*}_{c}\sim\gamma^{-1.8}$. Following the fitting results, we estimate the $l^{*}_{c}$ values for $\gamma\leq\gamma_{c}$, which is represented by the red star points. Then we compare these $l^{*}_{c}$ with actual simulated data for system size $1004$, and we get only a very mild signature of bending for $\gamma=0.04$ (which is still in the regime of the critical $\gamma_{c}$ with our estimated error-bar), but for $\gamma=0.03,0.02$ (they are $<\gamma_{c}$) the bending is not at all clearly evident for the mentioned system size. We can interpret this in two different ways, either $l^{*}_{c}$ scales faster than power-law ${}\gamma^{-1.8}$ in the case of small $\gamma$ limit, so that the $l^{*}_{c}$ (beyond which $c^{eff}_{l}$ will show the bending) is very large for small $\gamma$ values, which is not accessible numerically, or there exist a critical phase for $\gamma<\gamma_{c}$, in that case data will never bend. Results in the absence of periodic drive are also presented in the Appendix. D for our protocol for comparison.

### v3

@@ I Introduction
Another class of entanglement phase transitions arises from the continuous monitoring of quantum systems. The stochastic Schrödinger equation typically governs these open quantum dynamics, and averaging over different realizations of the states leads to the well-known Lindblad master equation dynamics Paule (2018). Since entanglement entropy is a nonlinear function of the state, calculating it for each trajectory separately and then averaging reveals a phase transition. In contrast, any linear function of the state is featureless and does not exhibit such transitions. Several studies have explored these entanglement phase transitions in both interacting and non-interacting systems, considering clean as well as disordered systems with long-range and short-range hopping, using both Lindblad jump operators and quantum state diffusion approaches Alberton et al. (2021); Szyniszewski et al. (2023); Cao et al. (2019); Minato et al. (2022a); Müller et al. (2022); Yamamoto and Hamazaki (2023); Fuji and Ashida (2020); Ladewig et al. (2022). Several studies have also been reported on the Ising chain with $Z_{2}$ symmetry under weak measurements accompanied by quantum jumps, demonstrating a transition between a critical phase and a Zeno phase Turkeshi et al. (2021, 2022); Le Gal et al. (2024). The impact of a single quantum jump on a localized site in a free fermionic system with $U(1)$ symmetry has been explored in Ref. Di Fresco et al. (2024), along with investigations into two-particle drive and loss processes in Ref. Soares et al. (2025). Additionally, the effect of different unravelings of the Lindblad master equation—namely, quantum state diffusion and quantum jumps—on the Ising chain is discussed in Ref. Piccitto et al. (2022). Interestingly, these two distinct measurement schemes lead to different patterns of entanglement growth. Further studies have examined the behavior of entanglement under long-range measured operators Piccitto et al. (2023); Russomanno et al. (2023), as well as the delocalization properties and bifurcation characteristics in the distribution of the expectations of the measured operators for a free fermionic system with $U(1)$ symmetry Piccitto et al. (2024). Continuously monitored open systems display competition between unitary evolution, which drives entanglement growth, and measurements, which tend to localize the wavefunction by collapsing it into eigenstates of the measurement operators. As a result, these systems undergo a measurement-induced phase transition as a function of measurement strength, which also has been reported recently in experiments Koh et al. (2023a). It has been shown that, in both clean and disordered free-fermionic systems, there is an entanglement phase transition from a critical to an area law phase as the strength of continuous measurements is tuned. This transition was predicted to fall into the Berezinskii-Kosterlitz-Thouless universality class Cao et al. (2019); Szyniszewski et al. (2023); Alberton et al. (2021), a result that was also supported by a theoretical study in which the dynamics of a continuously monitored free fermionic system were described by an effective non-Hermitian sine-Gordon model Buchhold et al. (2021). However, it is important to point out that the stability of the critical phase in a one-dimensional free-fermionic system has been questioned recently, and it has been argued in the thermodynamic limit, for any finite measurement strength, only the area-law phase will survive Poboiko et al. (2023); Starchl et al. (2024a); Poboiko et al. (2025) for random projective and quantum jump measurement protocols. Also in the case of Dirac fermions, the same conclusion is claimed recentlyMüller et al. (2025) after refining their previous resultsBuchhold et al. (2021). There are recent theoretical studies of a monitored chain of a system of size $L$ of Majorana fermions, which shows ($\ln L)^{2}$ behavior in the critical phase Fava et al. (2023), in contracts for noninteracting systems where the $U(1)$ symmetry is preserved, reported to show area-law phase for all measurement strength in presence of complex random hoppingFava et al. (2024). On the other hand, in the case of interacting systems, a transition from a volume law to an area law has been observed Fuji and Ashida (2020); Yamamoto and Hamazaki (2023); Xing et al. (2024). Furthermore, the role of measurements in localizing quantum states in nonunitary quantum circuits has also attracted significant interest Zabalo et al. (2022); Sierant and Turkeshi (2022); Oshima and Fuji (2023); Weinstein et al. (2023); Côté and Kourtis (2022); Block et al. (2022); Sang and Hsieh (2021); Lunt et al. (2021); Li et al. (2021); Fan et al. (2021); Zabalo et al. (2020); Turkeshi et al. (2020); Gullans and Huse (2020a, b); Bao et al. (2020); Choi et al. (2020); Tang and Zhu (2020); Jian et al. (2020). Some of these studies have also been extended in dimension $D>1$  Chahine and Buchhold (2024); Jin and Martin (2024); Poboiko et al. (2024a) and a semi-classical limit of such transition has also been introduced recently Ruidas and Banerjee (2024). In the case of $D>1$, Ref. Chahine and Buchhold (2024) maps a two-dimensional free fermionic monitored lattice system to a nonlinear sigma model (NLSM), suggesting that a $D$-dimensional free fermionic measurement problem can be mapped to a $(D+1)$-dimensional Anderson localization problem. If this mapping holds in one dimension, it implies that there would be no measurement-induced phase transition in the quantum state diffusion (QSD) protocol used in this manuscript. Additionally, for the two-dimensional bipartite lattice, the authors also map the monitored system to a non-Hermitian Hubbard model and observe numerical results that deviate from the predictions of the NLSM-based Anderson localization model. Specifically, they report the emergence of a conformal field theory (CFT) at the critical point—behavior that is absent in the Anderson localization scenario described by the NLSM. Since the Hubbard model mapping is valid for the continuous QSD protocol, the study raises important questions about the correspondence between the projective and QSD protocols, and about the broader validity of mapping a $D$-dimensional measurement problem to a $(D+1)$-dimensional Anderson localization problem. Notably, the projective measurement protocol shows strong parallels with the $(D+1)$-dimensional Anderson model applied to a $D$-dimensional free fermionic lattice system Poboiko et al. (2024b). In contrast to disorder-induced localization observed in the Anderson-localized or MBL phase, the measurement-induced localization is governed by a completely different mechanism and driven by quantum measurements and measurement strength.
@@
In parallel, in the last few decades, periodically driven (Floquet) quantum systems have emerged as a powerful framework for studying non-equilibrium phases of matter Banerjee and Sengupta (2024); Sarkar et al. (2021, 2022); Nag et al. (2014); Agarwala et al. (2016); Luitz et al. (2018); Chakrabarty and Datta (2024a). Floquet systems are known to exhibit exotic phases, such as discrete-time crystals Khemani et al. (2016); Else et al. (2016); Zhang et al. (2017) and Floquet topological insulators Oka and Aoki (2009); Lindner et al. (2011); Kitagawa et al. (2011); Kundu et al. (2014); Mukherjee et al. (2016). The fate of many-body localization (MBL) in the presence of a periodic drive has also been studied Lazarides et al. (2015); Ponte et al. (2015). While it has been demonstrated that Floquet systems can stabilize or even induce new phases, it remains an open question how continuous monitoring affects these driven systems and whether the introduction of a periodic drive can significantly alter the dynamics of a continuously monitored system, either enhancing or suppressing the effects of measurements. The effect of time-dependent Hamiltonian has been studied within the Lindblad master equation formalism Scopa et al. (2019); Ikeda et al. (2021), but the effect of periodic drive on a continuously monitored open quantum system, particularly in terms of trajectory-averaged entanglement entropy, remains relatively unexploredLi et al. (2019). In our study, we investigate such cases, and our main finding, based on our accessible system size numerics, is that periodically driven non-interacting systems tend to exhibit a signature of undergoing a BKT transition for reasonably high driving periods. In general, for a given drive protocol, the low-frequency regime favors the critical phase more, implying that if the BKT transition persists, then the critical measurement strength (transition point) increases with the time period. Moreover, the symmetric drive of the hopping amplitude about a zero mean is special; such a system always favors the area-law phase in the thermodynamic limit, independent of the frequency regime, in contrast to what has been observed in the case of the disorder-induced scenario Sarkar et al. (2021). We summarize our setup and main results based on the finite-sized numerical data in the form of a schematic in Fig. 1, which includes real data. We plot the maximum of the local effective central charge, which is a measure of the maximum growth rate of the steady state entanglement with sub-system size (defined later in the manuscript in Eq. (19)) with time-period $T$ for different measurement strength $\gamma$. It is increasingly believed, based on recent analytical studies, that in the absence of drive, even an infinitesimal measurement strength is sufficient to stabilize the area-law entanglement phase in the thermodynamic limit. In particular, analyses using the nonlinear sigma model (NLSM) have shown that the measurement-induced phase transition in one-dimensional free fermionic lattice models is bounded by a characteristic length scale, which grows exponentially as $\gamma$ decreases. However, clear numerical evidence remains lacking—especially for the specific measurement protocols considered in this work—due to the exponentially diverging correlation length. That said, for quantum jump and projective measurement protocols, numerical signatures consistent with area-law entanglement at small $\gamma$ have been reported Starchl et al. (2024b); Poboiko et al. (2023). Moreover, to the best of our knowledge, no analytical prediction is available for driven systems till now. This work aims to do a detailed comparative numerical analysis between the driven system with a reasonably low-frequency regime (in the main text) and the no-drive/high-frequency regime (in the Appendix D) results for a finite-sized system.
@@ IV.1 High-frequency regime: Analytical prediction
Note that this analysis has been questioned recently with the growing belief of the absence of measurement-induced phase transition in these models, and the analysis predicts a critical to area law phase transition in the no-drive scenario. However, since this study still captures the essential features of the measurement-induced transition at least in the case of a finite-size system, and can be extended to the driven system—at least in the extreme high-frequency limit—it motivates us to carry out this analysis to develop some intuition about the driven case.
@@
It is straightforward to check that for a fixed non-zero $\gamma$, in the limit $\epsilon\to 0$, this correlation tends to zero, and the effective central charge $c(\gamma)$ (note that the leading order term in the entanglement entropy $S=\frac{c(\gamma)}{3}\ln l$) will also approach zero as $\sqrt{\frac{\epsilon}{\gamma}}$. This implies that in the case of $\epsilon=0$, for any finite $\gamma$, one would expect to observe the area-law phase. In a nutshell in case of finite size systems, we make four main predictions based on this RG analysis about the high-frequency limit drive: 1) the entanglement dynamics can be mimicked by the dynamics of no-drive continuously monitored free-fermionic chain with the hopping amplitude $J_{eff}=\epsilon/2$, 2) If one starts in critical-phase in the no-drive regime, i.e. $\epsilon=2J$, with decreasing $\epsilon$, one can be able to probe critical to area-law phase transition, 3) $\epsilon=0$ limit will not display entanglement transition, it will always show the area-law entanglement for any non-zero finite measurement strength, 4) the transition point $\gamma_{c}$ is expected to decay to zero linearly with decreasing $\epsilon$ in the high-frequency region.
In this context, we would like to mention that very recently, a modified version of the RG analysis on sine-Gordon type models has been proposed after refining their previous analysis, in Ref. Müller et al. (2025), which shows the critical to area law transition point shifts towards zero in the case of a non-interacting monitored system in the absence of drive. However, the symmetries of these continuum models are not exactly identical to our lattice setup. In monitored systems, energy conservation is broken, and measurements can populate fermions across the entire band. Hence, there is no guarantee that any such models (that include Hamiltonian Eq. (9) too) can describe the dynamics of monitored fermions on a tight-binding lattice. Also, the prediction of Ref. Müller et al. (2025) of a logarithmic to area law phase transition in the interacting case, contradicts the other recent works Poboiko et al. (2025); Xing et al. (2024); Guo et al. (2024), where they observed a volume-law to area-law phase transition with increasing measurement rate. In our study, the RG scheme we used has helped us build an intuition about measurement-induced phase transitions in the presence of drive, which we will test in the subsequent section with finite-sized numerics.
@@ IV.2 Numerical results
As a beginning, we study the time evolution of the entanglement entropy, i.e., averaged over many different trajectories, starting from the initial zero entanglement product state for different values of $\epsilon$ and measurement strengths $\gamma,$ for $T=5$. Figure. 6 shows the dynamics of $\overline{S(nT)}$ for $\epsilon=3J/4$. The upper panel shows that for $\gamma=0.05$, the saturation value of the entanglement entropy increases with system size; on the other hand, the lower panel shows that for $\gamma=0.2$, the steady state value of the average entanglement entropy hardly changes with system size. This is a signature that, with increasing measurement strength, an entanglement transition may occur, where for small values of $\gamma$, the steady-state entanglement increases with $L$, and for large $\gamma$, it does not scale with $L$. The inset of the upper panel shows the entanglement dynamics in the high-frequency regime $T=0.5$ for $\gamma=0.05$ and $L=250$ (red-dotted line). As predicted in the previous section, we find that, indeed, for $T=0.5$, the entanglement dynamics can be mimicked by the continuously monitored no-drive scenario in a tight-binding lattice with an effective hopping strength $J_{eff}=\epsilon/2$, which is denoted by points in the inset figure. Given, the results for $T<1$ are almost indistinguishable from the no-drive results, at least for our accessible system size, hence, we consider $T=5>1$ for most of our numerics. The no-drive results are separately presented in the Appendix. D
@@
where, $\tilde{l}=\frac{L}{\pi}\sin(\pi l/L)$ is the cord length and $l$ is the size of the subsystem varies from $1$ to $L/2$. For logarithmic growth of the entanglement entropy, this measure should be constant; for volume-law scaling, it should scale as $c^{eff}_{l}\sim l$, while area-law behavior leads to $c^{eff}_{l}=0$. To calculate $c^{eff}_{l}$ numerically, we choose $N_{l}$ values of $l_{i}$ with $l_{1}=1$ and $l_{N_{l}}=L/2$ such that the successive ratio of $\frac{\tilde{l^{\prime}}}{\tilde{l}}$ is approximately constant. In Ref. Starchl et al. (2024b), it was shown that the critical phase is bounded from below by $l\sim\gamma^{-1}$ and from above by $l\sim\gamma^{-2}$. If the effective central charge increases with system size and attains a maximum, then it slowly starts decreasing, which indicates a crossover to area law. In Fig. 10 we present the effective central charge $c^{eff}_{l}$ data for the driving period $T=5$. We see a very clear evidence of bending of $c^{eff}_{l}$ after a critical $l^{*}_{c}$ for $\gamma>\gamma_{c}=0.05\pm 0.014$ (black circles represent the maximum). We fit those $l^{*}_{c}$ and obtain the scaling $l^{*}_{c}\sim\gamma^{-1.8}$. Following the fitting results, we estimate the $l^{*}_{c}$ values for $\gamma\leq\gamma_{c}$, which is represented by the red star points. Then we compare these $l^{*}_{c}$ with actual simulated data for system size $1004$, and we get only a very mild signature of bending for $\gamma=0.04$ (which is still in the regime of the critical $\gamma_{c}$ with our estimated error-bar), but for $\gamma=0.03,0.02$ (they are $<\gamma_{c}$) the bending is not at all clearly evident for the mentioned system size. We can interpret this in two different ways, either $l^{*}_{c}$ scales faster than power-law ${}\gamma^{-1.8}$ in the case of small $\gamma$ limit, so that the $l^{*}_{c}$ (beyond which $c^{eff}_{l}$ will show the bending) is very large for small $\gamma$ values, which is not accessible numerically, or there exist a critical phase for $\gamma<\gamma_{c}$, in that case data will never bend. Results in the absence of periodic drive are also presented in the Appendix. D for our protocol for comparison. We have also presented the equal time connected correlation function calculations in Appendix G as another tool to support the entanglement entropy data for $T=5$ for finite size systems.